OpenStudy (anonymous):

sq rt of -19 in terms of i?

OpenStudy (austinl):

What is the square root of 19? \(\sqrt{19}=~?\)

OpenStudy (austinl):

In exact terms mind you.

OpenStudy (anonymous):

\[\sqrt{-19} =\sqrt{-1 \times 19}=i \sqrt{19}\]

OpenStudy (anonymous):

so you have \[\sqrt{-19}\] remember that i is an imaginary number so it becomes \[i \sqrt{19}\]

OpenStudy (austinl):

Good job on giving the answer guys...